# From Restless to Contextual: A Thresholding Bandit Approach to Improve Finite-horizon Performance

### v2

@@ Appendix C Technical Details In Section 3
Therefore, the only thing we need to show is the Eq. (C.42) for $t=h$ under the assumption that Eq. (C.43) holds for $t=h$ and that Eq. (C.42) holds for $t=h+1$. We have
@@
$\forall t\leq T,$ $\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (C.44)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (C.45)
$\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (C.46)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (C.47)
$\pi_{t}^{\text{greedy}}=\pi_{t}^{*}.$ (C.48)
When $t=T$, this holds trivially, suppose this holds for $t\geq h+1$. When $t=h$, we will first show that Eq. (C.48) holds for $t=h$ under the assumption that Eq. (C.44), Eq. (C.45), Eq. (C.46) and Eq. (C.47) hold for $t=h+1$. This again is equivalent to show that
@@
×$\times\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-h}\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h}\right).$ (C.49)
@@
$\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h}\right)/d_{\text{upper}}$. Then to show that Eq. (C.49) is larger than 0, we only need to show that
@@
Next, we will first prove Eq. (C.44) for $t=h$ under the assumption that Eq. (C.44), Eq. (C.45), Eq. (C.46) and Eq. (C.47) hold for $t=h+1$ and Eq. (C.43) holds for $t=h$ We have
@@
Under the same condition, for Eq. (C.45) when $t=h$, we have
@@
where the inequality holds because we have $q^{1}_{11}-q^{1}_{01}\leq q^{0}_{11}-q^{0}_{01}$. Using the same logic will give us the Eq. (C.47) and Eq. (C.46) for $t=h$.
@@
$\mathcal{S}^{M-2}_{1}$ :=$:=\{\bm{s}:\bm{s}\in\mathcal{S}^{M-2},n_{\bm{s}}\geq B\}$
$\mathcal{S}^{M-2}_{2}$ :=$:=\{\bm{s}:\bm{s}\in\mathcal{S}^{M-2},n_{\bm{s}}=B-1\}$
$\mathcal{S}^{M-2}_{3}$ :=$:=\{\bm{s}:\bm{s}\in\mathcal{S}^{M-2},n_{\bm{s}}\leq B-2\},$
where $n_{\bm{s}}$ denotes the number of state 00 of $\bm{s}$.
@@
$\forall t\leq T,\forall\bm{s}\in\mathcal{S}^{M-2}_{3},$ $0\leq V_{t}^{*}((1,0,\bm{s}))-V_{t}^{*}((0,0,\bm{s}))\leq R(1)-R(0)$ (C.50)
$0\leq V_{t}^{*}((1,1,\bm{s}))-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (C.51)
$\forall\bm{s}\in\mathcal{S}^{M-2}_{2},$ $0\leq V_{t}^{*}((1,0,\bm{s}))-V_{t}^{*}((0,0,\bm{s}))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (C.52)
$0\leq V_{t}^{*}((1,1,\bm{s}))-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (C.53)
$\forall\bm{s}\in\mathcal{S}^{M-2}_{1},$ $0\leq V_{t}^{*}((1,0,\bm{s}))-V_{t}^{*}((0,0,\bm{s}))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (C.54)
$0\leq V_{t}^{*}((1,1,\bm{s}))-V_{t}^{*}((1,0,\bm{s}))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (C.55)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.56)
This holds trivially when $t=T$. Suppose $t\geq h+1$, this holds, we will first prove Eq. (C.56) for $t=h$ under the assumption that Eq. (C.50), Eq. (C.51), Eq. (C.52), Eq. (C.53), Eq. (C.54) and Eq. (C.55) hold for $t=h+1$. Similarly we have the following Bellman optimality equation.
@@
$\geq\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{1}_{1s_{h+1}^{1}}q^{0}_{0s_{h+1}^{2}}\sum_{\bm{s}_{h+1}}q_{\bm{s}_{h},\bm{s}_{h+1}}^{\bm{a}_{h+1}}V_{h+1}^{*}\left((s_{h+1}^{1},s_{h+1}^{2},\bm{s}_{h+1})\right).$ (C.57)
Suppose Eq. (C.57) holds, starting from any policy of allocating action that is not greedy, we can find an agent who is in state 1 and gets action 1 while another agent who is in state 0 gets action 0. By using the result, switching the action between them will give us a better policy. Suppose there are less than $B$ agents who are in state 0, then we can keep switching until all agents who are in state 0 are given action 1 , this is the greedy policy. Suppose there are more than $B$ agents who are in state 0, then we can keep switching until $B$ agents who are in state 0 get action 1, which is the greedy policy.
Now, we need to prove Eq. (C.57). Eq. (C.57) is equivalent to
@@
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,1)\right)-2F_{\bm{a}_{h+1},h+1}\left((1,0)\right)+F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)\geq 0,$ (C.58)
where $F_{\bm{a}_{h+1},h+1}\left((s_{h+1}^{1},s_{h+1}^{2})\right)=\sum_{\bm{s}_{h+1}}q_{\bm{s}_{h},\bm{s}_{h+1}}^{\bm{a}_{h+1}}V_{h+1}^{*}\left((s_{h+1}^{1},s_{h+1}^{2},\bm{s}_{h+1})\right)$. The reason why this holds is similar to what we proved in binary agent case. We know from the assumption that Eq. (C.50)-(C.55) hold and the definition of $F$ that
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,1))-F_{\bm{a}_{h+1},h+1}((1,0))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,0))-F_{\bm{a}_{h+1},h+1}((0,0))\leq\frac{R(1)-R(0)}{q^{0}_{10}}.$
@@
This shows Eq. (C.57), which, by our argument, shows Eq. (C.56) for $t=h$. Next, we will first define the specific greedy policy instead of randomly choosing agents who have higher incremental reward at time $h$ for different $\bm{s}$. When $\bm{s}\in\mathcal{S}^{M-2}_{1}$, we will only give action 1 to agent $3,4,\cdots,M$; when $\bm{s}\in\mathcal{S}^{M-2}_{2}$, we will give one action 1 to agent $1,2$; when $\bm{s}\in\mathcal{S}^{M-2}_{3}$, we will give two action 1 to agent $1,2$.
For $\bm{s}\in\mathcal{S}_{1}^{M-2}$, we will show Eq. (C.54) holds for $t=h$ under the assumption that Eq.(C.50)-(C.55) hold for $t=h+1$ and Eq. (C.56) holds for $t=h$. We have
@@
Next, we will show the RHS of Eq. (C.54) for $t=h$ under the same condition.
@@
Meanwhile, when $t=h$, again under the same condition, the LHS of Eq. (C.55) satisfies
@@
With the same reasoning, we can show the RHS of Eq. (C.55) under the same condition for $t=h$.
@@
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))\leq R(1)-R(0)$ (C.59)
$0\leq V_{t}^{*}\left((1,1,\bm{s})\right)-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (C.60)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.61)
They hold trivially when $t=T$. Suppose they hold when $t\geq h+1$, when $t=h$, we will first prove Eq. (C.61) for $t=h$ under the assumption that Eq. (C.59) and Eq. (C.38) hold for $t=h+1$. Using the same reasoning, we only need to show Eq. (C.58). Again, from the assumptions and the definition of $F$, we know that
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,1))-F_{\bm{a}_{h+1},h+1}((1,0))\leq{R(1)-R(0)}$
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,0))-F_{\bm{a}_{h+1},h+1}((0,0))\leq{R(1)-R(0)}.$

### v4

@@ Appendix E Technical Details In Section 3
Therefore, the only thing we need to show is the Eq. (E.28) for $t=h$ under the assumption that Eq. (E.29) holds for $t=h$ and that Eq. (E.28) holds for $t=h+1$. We have
@@
$\forall t\leq T,$ $\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (E.30)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (E.31)
$\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (E.32)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (E.33)
$\pi_{t}^{\text{greedy}}=\pi_{t}^{*}.$ (E.34)
When $t=T$, this holds trivially, suppose this holds for $t\geq h+1$. When $t=h$, we will first show that Eq. (E.34) holds for $t=h$ under the assumption that Eq. (E.30), Eq. (E.31), Eq. (E.32) and Eq. (E.33) hold for $t=h+1$. This again is equivalent to show that
@@
$\times\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-h}\right)$
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h}\right).$ (E.35)
@@
$\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-h}\right)/d_{\text{upper}}$. Then to show that Eq. (E.35) is larger than 0, we only need to show that
@@
Next, we will first prove Eq. (E.30) for $t=h$ under the assumption that Eq. (E.30), Eq. (E.31), Eq. (E.32) and Eq. (E.33) hold for $t=h+1$ and Eq. (E.29) holds for $t=h$ . We have
@@
Under the same condition, for Eq. (E.31) when $t=h$, we have
@@
where the inequality holds because we have $q^{1}_{11}-q^{1}_{01}\leq q^{0}_{11}-q^{0}_{01}$. Using the same logic will give us the Eq. (E.33) and Eq. (E.32) for $t=h$.
Case 4:
Suppose $q^{1}_{11}\leq q^{1}_{01},q_{11}^{0}\geq q_{01}^{0},q_{11}^{0}\geq q_{11}^{1}$. We will directly show that $V_{t}^{\bm{\pi}^{\text{greedy}}}$ satisfies the optimal Bellman equation. With the same logic, we only need to prove that Eq. (E.22) holds for each $t\leq T$. We will first assume that the following holds and will prove this later:
$V_{t}^{\bm{\pi}^{\text{greedy}}}((1,1))+V_{t}^{\bm{\pi}^{\text{greedy}}}((0,0))-2V_{t}^{\bm{\pi}^{\text{greedy}}}((1,0))\leq 0,~\forall t\leq T.$ (E.36)
Then we have
$\left(q_{11}^{0}+q_{01}^{1}-q_{11}^{1}-q_{01}^{0}\right)\left(V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))-V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((0,0))\right)$
$+\left(q_{01}^{1}q_{11}^{0}-q_{01}^{0}q_{11}^{1}\right)\left(V_{t+1}^{\bm{\pi}^{\text{greedy}}}((1,1))+V_{t+1}^{\bm{\pi}^{\text{greedy}}}((0,0))-2V_{t+1}^{\bm{\pi}^{\text{greedy}}}((1,0))\right)$
$=\left(q_{11}^{0}+q_{01}^{1}-q_{11}^{1}-q_{01}^{0}\right)\left(V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))-V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((0,0))\right)$
$+\left(q_{11}^{1}(q_{11}^{0}-q_{01}^{0})-q_{11}^{0}(q_{11}^{1}-q_{01}^{1})\right)\left(V_{t+1}^{\bm{\pi}^{\text{greedy}}}((1,1))+V_{t+1}^{\bm{\pi}^{\text{greedy}}}((0,0))-2V_{t+1}^{\bm{\pi}^{\text{greedy}}}((1,0))\right)$
$\geq\left(q_{11}^{0}+q_{01}^{1}-q_{11}^{1}-q_{01}^{0}\right)\left(V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))-V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((0,0))\right)$
$+\left(q_{01}^{1}(q_{11}^{0}-q_{01}^{0})-q_{11}^{0}(q_{11}^{1}-q_{01}^{1})\right)\left(V_{t+1}^{\bm{\pi}^{\text{greedy}}}((1,1))+V_{t+1}^{\bm{\pi}^{\text{greedy}}}((0,0))-2V_{t+1}^{\bm{\pi}^{\text{greedy}}}((1,0))\right)$
$=-V_{t}^{\bm{\pi}^{\text{greedy}}}((1,1))-V_{t}^{\bm{\pi}^{\text{greedy}}}((0,0))+2V_{t}^{\bm{\pi}^{\text{greedy}}}((1,0))$
$\geq 0.$
Next, we will prove Eq. (E.36). To ease notation, we will let $a:=1-q_{01}^{1},b:=1-q_{11}^{0},c=q_{11}^{0}-q_{01}^{0},d=q_{11}^{1}-q_{01}^{1}$. By Eq. (E.13) and (E.14), we have the following iterative equations
$V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))-V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((0,0))$ $=R(1)-R(0)+ac\left(V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))-V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((0,0))\right)$
$+(1-a)c\left(V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((1,1))-V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))\right)$
$V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,1))-V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))$ $=R(1)-R(0)+bd\left(V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))-V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((0,0))\right)$
$+(1-b)d\left(V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((1,1))-V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))\right).$
Let $\bm{V}_{t}=\begin{pmatrix}V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))-V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((0,0))\\ V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,1))-V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))\end{pmatrix},A=\begin{pmatrix}ac&(1-a)c\\ bd&(1-b)d\end{pmatrix}$. We have
$\bm{V}_{t}=(R(1)-R(0))\bm{1}+A\bm{V}_{t+1}.$
First we consider the subcase where IAI-A is not invertible, which is $A$ has eigenvalue 1. This is equivalent to $\boxed{q_{11}^{0}=1,q_{01}^{0}=q_{01}^{1}=q_{11}^{1}=0}$. Under this case, we have $V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,1))-V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))=R(1)-R(0)$. We could also easily use induction to show that
$V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))-V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((0,0))=R(1)-R(0)+V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0))-V_{t+1}^{\bm{\pi}^{\text{\text{greedy}}}}((0,0))\geq R(1)-R(0).$
This will give us the desired result.
Next, we assume IAI-A is invertible, then we have
$\bm{V}_{t}$ $=(R(1)-R(0))(I-A)^{-1}\bm{1}+(R(1)-R(0))A^{T-t}\left(\bm{1}-(I-A)^{-1}\bm{1}\right)$
$=(R(1)-R(0))(I-A)^{-1}\bm{1}-(R(1)-R(0))A^{T-t+1}(I-A)^{-1}\bm{1}$
$=\sum_{k=0}^{T-t}(R(1)-R(0))A^{k}\bm{1}.$
It is easy to see that $A$ has two eigenvalues:
$\lambda_{1}=\frac{\eta+\sqrt{\eta^{2}-4\delta}}{2}~,\lambda_{2}=\frac{\eta-\sqrt{\eta^{2}-4\delta}}{2},$
where $\eta=ac+(1-b)d,\delta=(a-b)cd$.
Then, we first consider the case $\boxed{\lambda_{i}\in\mathbb{R}}$. We have
$-V_{t}^{\bm{\pi}^{\text{greedy}}}((1,1))-V_{t}^{\bm{\pi}^{\text{greedy}}}((0,0))+2V_{t}^{\bm{\pi}^{\text{greedy}}}((1,0))$ $=(1,-1)\bm{V}_{t}\bm{1}$
$=(R(1)-R(0))(c-d)\sum_{k=0}^{T-t}\frac{\lambda_{1}^{k}-\lambda_{2}^{k}}{\lambda_{1}-\lambda_{2}}.$
Since under Case 4, $c\geq d$, we only need to prove that $\sum\frac{\lambda_{1}^{k}-\lambda_{2}^{k}}{\lambda_{1}-\lambda_{2}}\geq 0$.
When $\eta\geq 0$, $|\lambda_{1}|\geq|\lambda_{2}|$. Therefore, for all $k$, $\lambda_{1}^{k}\geq\lambda_{2}^{k}$, which will give us the desired result.
When $\eta\leq 0,\delta\geq 0$, we have $1\leq\eta\leq\lambda_{2}\leq\lambda_{1}\leq 0$. First, we will prove $\lambda_{1}+\lambda_{1}^{2}\geq\lambda_{2}+\lambda_{2}^{2}$:
$\lambda_{1}+\lambda_{1}^{2}-\lambda_{2}-\lambda_{2}^{2}$ $=(\lambda_{1}-\lambda_{2})(\lambda_{1}+\lambda_{2}+1)$
$=\delta(1+\eta)$
$\geq 0.$
Therefore, we have $\frac{\lambda_{1}}{\lambda_{2}}\leq\frac{1+\lambda_{2}}{1+\lambda_{1}}$. Then we can have for each odd $k$
$\frac{\lambda_{1}^{k}}{\lambda_{2}^{k}}\leq\frac{\lambda_{1}}{\lambda_{2}}\leq\frac{1+\lambda_{2}}{1+\lambda_{1}}.$
This means for each odd $k$, $\lambda_{1}^{k}+\lambda_{1}^{k+1}\geq\lambda_{2}^{k}+\lambda_{2}^{k+1}$. Then we have for all $T-t,\sum_{k=0}^{T-t}\lambda_{1}^{k}-\lambda_{2}^{k}\geq 0$. This will give us the desired result.
When $\eta\leq 0,\delta\leq 0$, we can numerically show that $\lambda_{2}\geq-1$. Specifically, we use gradient descent with a sufficient small step size ($10^{-5}$) and multiple initialization by choosing the grid points to calculate the solutions of the KKT conditions. The results shows that the minimum is larger than 1-1. Then we have $\sum_{k=0}^{T-t}\lambda_{2}^{k}\leq 1$. We also have that $\sum_{k=0}^{T-t}\lambda_{1}^{k}\geq 1$ since $\lambda_{1}\geq 0$. This will give us the desired result.
Finally, we consider the subcase where $\boxed{\lambda_{i}\in\mathbb{C}}$. We note that this will only happen when $\delta\geq 0$ which means $q_{11}^{0}\leq q_{01}^{1}$. We will first rewrite the eigenvalues using its spherical form, i.e. $\lambda_{i}=\rho\exp(\pm i\theta)$, where
$\rho=\sqrt{(a-b)cd}~,\cos(\theta)=\frac{ac+(1-b)d}{2\sqrt{(a-b)cd}}.$
Then we have
$-V_{t}^{\bm{\pi}^{\text{greedy}}}((1,1))-V_{t}^{\bm{\pi}^{\text{greedy}}}((0,0))+2V_{t}^{\bm{\pi}^{\text{greedy}}}((1,0))=(1,-1)\bm{V}_{t}\bm{1}$
$=(R(1)-R(0))(c-d)\sum_{k=0}^{T-t}\frac{\lambda_{1}^{k}-\lambda_{2}^{k}}{\lambda_{1}-\lambda_{2}}$
$=(R(1)-R(0))\frac{c-d}{1-2\rho\cos(\theta)+\rho^{2}}\left(1-\rho^{T-t}\frac{\sin((T-t+1)\theta-\rho\sin((T-t)\theta))}{\sin(\theta)}\right).$
Since $c-d\geq 0,1-2\rho\cos(\theta)+\rho^{2}=|1-\rho\exp(i\theta)|^{2}\geq 0$, we only need to show that
$1-\rho^{T-t}\frac{\sin((T-t+1)\theta-\rho\sin((T-t)\theta))}{\sin(\theta)}\geq 0.$
Since $0\leq\rho\leq 1$, we only need to show the following:
$\rho^{T-t}\frac{\sin((T-t+1)\theta-\sin((T-t)\theta))}{\sin(\theta)}$ $\leq 1$
$\rho^{T-t}\frac{\sin((T-t+1)\theta)}{\sin(\theta)}$ $\leq 1.$
We will use induction to prove the following:
$\rho^{k}\frac{\sin((k+1)\theta-\sin((k)\theta))}{\sin(\theta)}$ $\leq 1$
$\rho^{k}\frac{\sin((k+1)\theta)}{\sin(\theta)}$ $\leq 1.$
When $k=0$, the statements trivially hold. Suppose it holds for $k=t$, when $k=t+1$, we have
$\rho^{t+1}\frac{\sin((t+2)\theta-\sin((t+1)\theta))}{\sin(\theta)}$ $=\rho^{t+1}\frac{\cos((t+1)\theta+\theta/2)}{\cos(\theta/2)}$
$=\rho^{t+1}\frac{\cos(t\theta+\theta/2)\cos(\theta)-\sin(t\theta+\theta/2)\sin(\theta)}{\cos(\theta/2)}$
$\leq\rho\cos(\theta)+2\sin(\theta/2)\rho.$
Similarly, we have
$\rho^{t+1}\frac{\sin((t+2)\theta)}{\sin(\theta)}$ $=\rho^{t+1}\frac{\sin((t+1)\theta)\cos(\theta)+\sin(\theta)\cos((t+1)\theta)}{\sin(\theta)}$
$\leq\rho\cos(\theta)+\rho$
$=\frac{1}{2}\left(ac+(1-b)d+2\sqrt{(q_{01}^{1}-q_{11}^{1})(q_{01}^{1}-q_{11}^{0})(q_{11}^{0}-q_{01}^{0})}\right)$
$\leq\frac{1}{2}\left((1-q_{01}^{1})q_{11}^{0}+q_{11}^{0}(q_{11}^{1}-q_{01}^{1})+2(q_{01}^{1}-q_{11}^{1})\right)$
$\leq\frac{1}{2}\left((1-q_{01}^{1})q_{11}^{0}+(2-q_{11}^{0})q_{01}^{1}\right)$
$\leq 1.$
@@
$\mathcal{S}^{M-2}_{1}$ $:=\{\bm{s}:\bm{s}\in\mathcal{S}^{M-2},n_{\bm{s}}\geq B\}$
$\mathcal{S}^{M-2}_{2}$ $:=\{\bm{s}:\bm{s}\in\mathcal{S}^{M-2},n_{\bm{s}}=B-1\}$
$\mathcal{S}^{M-2}_{3}$ $:=\{\bm{s}:\bm{s}\in\mathcal{S}^{M-2},n_{\bm{s}}\leq B-2\},$
where $n_{\bm{s}}$ denotes the number of state 0 of $\bm{s}$.
@@
$\forall t\leq T,\forall\bm{s}\in\mathcal{S}^{M-2}_{3},$ $0\leq V_{t}^{*}((1,0,\bm{s}))-V_{t}^{*}((0,0,\bm{s}))\leq R(1)-R(0)$ (E.37)
$0\leq V_{t}^{*}((1,1,\bm{s}))-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (E.38)
$\forall\bm{s}\in\mathcal{S}^{M-2}_{2},$ $0\leq V_{t}^{*}((1,0,\bm{s}))-V_{t}^{*}((0,0,\bm{s}))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (E.39)
$0\leq V_{t}^{*}((1,1,\bm{s}))-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (E.40)
$\forall\bm{s}\in\mathcal{S}^{M-2}_{1},$ $0\leq V_{t}^{*}((1,0,\bm{s}))-V_{t}^{*}((0,0,\bm{s}))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (E.41)
$0\leq V_{t}^{*}((1,1,\bm{s}))-V_{t}^{*}((1,0,\bm{s}))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (E.42)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (E.43)
This holds trivially when $t=T$. Suppose $t\geq h+1$, this holds, we will first prove Eq. (E.43) for $t=h$ under the assumption that Eq. (E.37), Eq. (E.38), Eq. (E.39), Eq. (E.40), Eq. (E.41) and Eq. (E.42) hold for $t=h+1$. Similarly we have the following Bellman optimality equation.
@@
$\geq\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{1}_{1s_{h+1}^{1}}q^{0}_{0s_{h+1}^{2}}\sum_{\bm{s}_{h+1}}q_{\bm{s}_{h},\bm{s}_{h+1}}^{\bm{a}_{h+1}}V_{h+1}^{*}\left((s_{h+1}^{1},s_{h+1}^{2},\bm{s}_{h+1})\right).$ (E.44)
Suppose Eq. (E.44) holds, starting from any policy of allocating action that is not greedy, we can find an agent who is in state 1 and gets action 1 while another agent who is in state 0 gets action 0. By using the result, switching the action between them will give us a better policy. Suppose there are less than $B$ agents who are in state 0, then we can keep switching until all agents who are in state 0 are given action 1 , this is the greedy policy. Suppose there are more than $B$ agents who are in state 0, then we can keep switching until $B$ agents who are in state 0 get action 1, which is the greedy policy.
Now, we need to prove Eq. (E.44). Eq. (E.44) is equivalent to
@@
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(F_{\bm{a}_{h+1},h+1}\left((1,1)\right)-2F_{\bm{a}_{h+1},h+1}\left((1,0)\right)+F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)\geq 0,$ (E.45)
where $F_{\bm{a}_{h+1},h+1}\left((s_{h+1}^{1},s_{h+1}^{2})\right)=\sum_{\bm{s}_{h+1}}q_{\bm{s}_{h},\bm{s}_{h+1}}^{\bm{a}_{h+1}}V_{h+1}^{*}\left((s_{h+1}^{1},s_{h+1}^{2},\bm{s}_{h+1})\right)$. The reason why this holds is similar to what we proved in binary agent case. We know from the assumption that Eq. (E.37)-(E.42) hold and the definition of $F$ that
$0$ $\leq F_{\bm{a}_{h+1},h+1}((1,1))-F_{\bm{a}_{h+1},h+1}((1,0))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$
$0$ $\leq F_{\bm{a}_{h+1},h+1}((1,0))-F_{\bm{a}_{h+1},h+1}((0,0))\leq\frac{R(1)-R(0)}{q^{0}_{10}}.$
@@
This shows Eq. (E.44), which, by our argument, shows Eq. (E.43) for $t=h$. Next, we will first define the specific greedy policy instead of randomly choosing agents who have higher incremental reward at time $h$ for different $\bm{s}$. When $\bm{s}\in\mathcal{S}^{M-2}_{1}$, we will only give action 1 to agent $3,4,\cdots,M$; when $\bm{s}\in\mathcal{S}^{M-2}_{2}$, we will give one action 1 to agent $1,2$; when $\bm{s}\in\mathcal{S}^{M-2}_{3}$, we will give two action 1 to agent $1,2$.
For $\bm{s}\in\mathcal{S}_{1}^{M-2}$, we will show Eq. (E.41) holds for $t=h$ under the assumption that Eq.(E.37)-(E.42) hold for $t=h+1$ and Eq. (E.43) holds for $t=h$. We have
@@
Next, we will show the RHS of Eq. (E.41) for $t=h$ under the same condition.
@@
Meanwhile, when $t=h$, again under the same condition, the LHS of Eq. (E.42) satisfies
@@
With the same reasoning, we can show the RHS of Eq. (E.42) under the same condition for $t=h$.
@@
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))\leq R(1)-R(0)$ (E.46)
$0\leq V_{t}^{*}\left((1,1,\bm{s})\right)-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (E.47)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (E.48)
They hold trivially when $t=T$. Suppose they hold when $t\geq h+1$, when $t=h$, we will first prove Eq. (E.48) for $t=h$ under the assumption that Eq. (E.46) and Eq. (E.24) hold for $t=h+1$. Using the same reasoning, we only need to show Eq. (E.45). Again, from the assumptions and the definition of $F$, we know that
$0$ $\leq F_{\bm{a}_{h+1},h+1}((1,1))-F_{\bm{a}_{h+1},h+1}((1,0))\leq{R(1)-R(0)}$
$0$ $\leq F_{\bm{a}_{h+1},h+1}((1,0))-F_{\bm{a}_{h+1},h+1}((0,0))\leq{R(1)-R(0)}.$